# Multi-wavelength Emission from Jets and Magnetically Arrested Disks in Nearby Radio Galaxies: Application to M87

### v1

@@ 2.4 Particle distribution inside the jet
and $\gamma_{\rm rad}$ is the Lorentz factor estimated by balancing the acceleration and cooling timescales. We estimate the reconnection acceleration timescale as $t_{\rm acc}=\eta_{\rm acc}\gamma m_{e}c/(eB_{\rm dis})$, [^3] If electrons are only accelerated by the reconnection electric field of the turbulent magnetic field, the acceleration timescale is estimated to be $t_{\rm acc}=\eta_{\rm acc}\gamma m_{e}c/(e\delta B)$. We estimate $\delta B$ as $\delta B\approx(\delta B/B)B_{\rm dis}$, and thus, the acceleration timescale is longer than that given in the text by a factor of $B/\delta B$. This affects the estimate of $\gamma_{\rm rad}$ but does not affect our results within the parameters of our interest. where $\eta_{\rm acc}\sim 10$ is the numerical factor (e.g., Guo et al., 2020). We estimate $B_{\rm dis}$ by Equation (12) as
@@
We consider the cooling processes of the electrons as synchrotron radiation, synchrotron self-Compton scattering, and adiabatic cooling. We estimate the cooling timescale as $t_{\rm cool}^{-1}=t_{\rm syn}^{-1}+t_{\rm SSC}^{-1}+t_{\rm ad}^{-1}$. We estimate the synchrotron cooling timescale to be $t_{\rm syn}=6\pi m_{e}c/(\gamma\sigma_{T}B_{\rm dis}^{2})$, where $\sigma_{T}$ is the Thomson cross section.
We calculate the synchrotron self-Compton scattering cooling timescale in the way of subsection 2.6 of Blumenthal & Gould (1970). We ignore the effect of the external photon field in this paper. The accretion rate is low for M87, and thus, the energy density of the external photon, such as thermal emission from a dust torus, is lower than that of the internal photon emitted by the energetic electrons.
As for the escape process, we consider that the electrons escape from the jet diffusely via resonant scattering with the turbulence. KHI injects the turbulence with the scale of the dissipation region, $R_{j}$, and cascades into the small scale. We assume that the spectrum of the turbulence is the classical Kolmogorov turbulence (e.g., San & Maulik, 2018). A turbulent eddy whose scale is the Larmor radius of the electrons, $r_{L}=\gamma m_{e}c^{2}/(eB_{\rm dis})$, scatters the electrons. We estimate the diffusion timescale as $t_{\rm esc}=R_{j}^{2}/D_{R}$, where $D_{R}\approx(4/\pi)r_{L}c(R_{j}/r_{L})^{2/3}(B/\delta B)^{2}$ is the diffusion coefficient (see also Kulsrud, 2005).
@@ 3 Results for strong plasma entrainment
In this section, we apply the Jet-MAD model to M87 and compare the multi-wavelength photon spectra obtained by the Jet-MAD model to the observational data. We tabulate the parameters and the physical quantities for M87 in Table 1. The other parameters in the MAD model are the same as those in Kimura & Toma (2020). We show the various timescales of the MAD model as a function of the proton energy $E_{p}$ for M87 in Figure 3. Due to the hard-sphere-like diffusion coefficient, the diffusion and acceleration timescales of the MAD model are independent of $E_{p}$. The cooling timescales are negligibly small for $E_{p}\lesssim 10^{9}~{}\rm GeV$.
The value of $s_{\rm inj}$ is given by the stochastic acceleration process with steady-state and hard-sphere approximations. As long as the cooling timescale is negligible, this process results in $s_{\rm inj}=-(1/2)+\sqrt{(9/4)+\epsilon}$, where $\epsilon=t_{\rm acc}/t_{\rm diff}$ is the escape-acceleration timescale ratio (see Section 2 of Stawarz & Petrosian, 2008). We obtain $s_{\rm inj}\simeq 1.1$ in our fiducial parameters.
@@
Parameters of the Jet-MAD model for strong plasma entrainment $\dot{m}$ $M_{\rm BH}[M_{\odot}]$ $r$ $s_{\rm{inj}}$ $Q_{i}/Q_{e}$ $\eta_{\rm turb}$ $\alpha$ $\beta$ $\sigma_{\rm ent}$ $p$ $\xi$ $\delta B/B$ $R_{j}[R_{g}]$ $6\times 10^{-5}$ $6.3\times 10^{9}$ 10 1.1 10 0.833 0.3 0.1 $0.01$ $2.16$ $10^{3}$ 0.35 $10^{2}$
Derived physical quantities of M87 in the jet $\sigma_{\pm}$ $B_{\rm dis}\rm[G]$ $n_{e}\rm\ [cm^{-3}]$ $L_{j}\ \rm[erg~{}s^{-1}]$ $L_{e}\ \rm[erg~{}s^{-1}]$ $\delta_{D}$ $1.0\times 10^{5}$ $6.5\times 10^{-2}$ $22$ $2.8\times 10^{43}$ $1.7\times 10^{41}$ 1.9
Figure 4: Top panel: Photon spectrum for M87. The thick and thin lines are the photon spectra after and before internal attenuation by the Breit-Wheeler process, respectively. Gray points are the multi-wavelength simultaneous data taken from EHT MWL Science Working Group et al. (2021) and MAGIC Collaboration et al. (2020). Blue and orange shaded regions are power-law and log-parabola fitting spectrum taken from the Fermi-4FGL-DR3 catalog (https://fermi.gsfc.nasa.gov/ssc/data/access/lat/12yr_catalog/). Bottom panel: Same as Figure 3, but timescales in the jet as a function of the electron energy. $t_{\rm SSC}$ (dotted) and $t_{\rm ad}$ (dot-dashed) are the synchrotron self-Compton and adiabatic cooling timescales, respectively. The vertical dashed gray line represents $E_{e,\rm cut}=\gamma_{\rm cut}m_{e}c^{2}$.
First, we consider a case with $(\delta B/B)^{2}\sigma_{\rm ent}<1$. We show the results in Figure 4. Thermal synchrotron radiation reproduces the submillimeter data. Synchrotron radiation by the nonthermal protons and secondary electron-positron pairs produced via the Bethe-Heitler process in the MAD reproduce the GeV and TeV gamma-ray data, respectively. Synchrotron radiation by the jet contributes to the optical, UV, and X-ray data. Our model can reproduce multi-wavelength data except for tens of GHz radio data.
@@
×$\times\xi^{2}_{3}\sigma^{2}_{\pm,4}\left(\frac{\sigma_{\rm dis}/(1+0.5\sigma_{\rm dis})}{0.66}\right)^{1/2}\ \rm eV.$ (18)
Thus, the electrons accelerated by the magnetic reconnection have a Lorentz factor high enough to emit the optical to the X-ray photons via synchrotron radiation. On the other hand, the photon flux of the synchrotron-self Compton scattering is much lower than the synchrotron flux as shown in the top panel of Figure 4. In our model, synchrotron flux is calibrated by the optical data to be $F_{\rm opt}\simeq 8\times 10^{-13}\ \rm erg~{}s^{-1}~{}cm^{-2}$, and we estimate the photon energy density as $U_{\rm rad}=F_{\rm opt}D^{2}/(R_{j}^{2}c)\simeq 8.4\times 10^{-6}\ \rm erg~{}cm^{-3}$, where $D=17~{}\rm Mpc$ is the distance to M87. We also estimate the magnetic energy density to be $U_{B}=B_{\rm dis}^{2}/(8\pi)\simeq 1.7\times 10^{-4}\ \rm erg~{}cm^{-3}$. The energy density ratio is estimated to be $U_{\rm rad}/U_{B}\simeq 0.05$, and thus the photon flux of the synchrotron-self Compton scattering is much lower than the synchrotron component.
The bottom panel of Figure 4 shows the various timescales in the jet as a function of the electron energy. For our fiducial parameters, electrons are efficiently scattered by the turbulence, and the diffusive escape is inefficient. The nonthermal electrons are mainly cooled by the adiabatic expansion and the synchrotron radiation for $\gamma<1.9\times 10^{4}$ and $\gamma>1.9\times 10^{4}$, respectively. The photon energy emitted by the electrons with $t_{\rm ad}=t_{\rm syn}$ is estimated to be $E_{\gamma,\rm cool}\simeq 0.29~{}\rm eV$, at which the synchrotron spectrum is the peak, and that enables our model to explain the optical data.
@@ 4 Resutlts for weak or no plasma entrainment
The obtained minimum Lorentz factor is too high to emit the optical and X-ray photons. Based on Equation (19), $\delta B/B\sim 0.01$ is required to explain the X-ray data. Even if we tune $\delta B/B$ to emit the X-ray photons by electrons with $\gamma_{\rm min}$, the photon spectrum is too peaky to fit the optical and the X-ray data simultaneously. We show the result in Figure 5 for the case of weak plasma entrainment. We conclude that the weak plasma entrainment is disfavored to explain the multi-wavelength data.
Figure 5: Same as Figure 4, but for the weak plasma entrainment. Total and jet synchrotron photon spectra with $\delta B/B=0.033$ (thick lines) and $\delta B/B=0.02$ (thin lines). We set $\delta B/B=0.02$ to emit the optical photons and $\delta B/B=0.033$ to emit the X-ray photons. Both photon spectra cannot reproduce the multi-wavelength data simultaneously.
@@ 5.1.1 Power-law tail in the jet dissipation
We apply the Jet-MAD model with the power-law tail distribution to M87 and find that the Jet-MAD model can explain the multi-wavelength data even if the differential electron number spectrum has the power-law tail. The index of synchrotron flux is $s\approx(2-p)/2$ for the fast cooling regime. We obtain $s=-0.25$ and $-0.5$ for $p=2.5$ and 3.0, respectively, with which the synchrotron flux emitted by the electrons in the tail ($\gamma>\gamma_{\rm cut}$) has little effect on the overall photon spectrum. Even if the tail index is a bit hard, the photon energy higher than $E_{\gamma,\rm max}$ does not affect the photon spectrum (see top panel of Figure 4). We conclude that our model can explain the multi-wavelength data independent of the uncertainty for the reconnection acceleration.
@@ 5.1.2 Power-law tail for the electron-positron pair injection into the jet
Next, we evaluate the effect of the power-law tail on the electron-positron pair injection at the jet base. With the power-law tail, the photon spectrum extends up to a higher energy than that without the tail. One might expect a high pair production efficiency, compared to that estimated with the exponential cutoff as in Chen et al. (2023). We estimate the pair production rate and magnetization parameter at the jet base with the power-law tail distribution (see Appendix A for details). We find that the estimated magnetization parameters with the tail are comparable to that in our fiducial treatment: $\sigma_{\pm}=8.8\times 10^{4}$ and $1.0\times 10^{5}$ for $p_{\rm tail}=2.5$ and $3.0$, respectively. We show the estimated magnetization parameter as a function of the tail index in Figure 6. As shown in Figure 6, the estimated magnetization parameter hardly changes its value with the tail index. This comes from the self-regulation of the specific luminosity to peak at $E_{\gamma,\rm cut,ms}\sim m_{e}c^{2}$. Therefore, the power-law tail does not affect our conclusions.
@@ 5.2 Acceleration by the Alfvén wave dissipation
In this subsection, we discuss the Alfvén waves dissipation as an alternative particle acceleration mechanism in the jet (Nättilä & Beloborodov, 2022). Alfvén waves are produced by the magnetic reconnection in the BH magnetosphere or by the interaction of the jet with disk wind. These Alfvén waves propagate along the magnetic field lines and split into forward sound waves and backward Alfvén waves by the parametric decay instability (PDI) (Galeev & Oraevskii, 1963; Sagdeev & Galeev, 1969; Derby, 1978). We consider that the Alfvén wave turbulence is generated by the interaction of the forward Alfvén waves with backward Alfvén waves at the dissipation region, and the dissipation of Alfvén wave turbulence accelerates the electrons. It is shown that in a highly magnetized medium, the PDI can split the Alfvén waves into backward Alfvén waves and forward sound waves only when $(\delta B/B)^{2}\sigma\leq 1$ is satisfied (Ishizaki & Ioka, 2024). We set $\delta B/B=1/\sqrt{\sigma_{\pm}}\simeq 3\times 10^{-3}$ which is close to the maximum value of $\delta B/B$ that satisfies the PDI growth condition. Then, the maximum Lorentz factor is written as $\gamma_{\rm max}\approx\xi$ by Equation (15), and the value is lower than that of the magnetic reconnection scenario.
We show the results in Figure 7. We estimate the synchrotron photon energy emitted by electrons with $\gamma_{\rm max}$ to be
@@
Diffusive escape also affects the photon spectrum. For the Alfvén wave dissipation scenario, the amplitude of the turbulent magnetic field is lower than that of the magnetic reconnection scenario. Due to the low amplitude, the pitch angle scattering rate is low, leading to a high spatial diffusion coefficient. The diffusive escape timescale is always shorter than the adiabatic and synchrotron cooling timescales, causing a low radiation efficiency, compared to the magnetic reconnection scenario (see bottom panel of Figure 7). As a result, the calculated photon flux is low for the Alfvén wave dissipation scenario. We conclude that the dissipation of the Alfvén wave turbulence is disfavored as the acceleration mechanism inside the jet.

### v2

@@ 2.4 Particle distribution inside the jet
and $\gamma_{\rm rad}$ is the Lorentz factor estimated by balancing the acceleration and cooling timescales. We estimate the reconnection acceleration timescale as $t_{\rm acc}=\eta_{\rm acc}\gamma m_{e}c/(eB_{\rm dis})$, [^2] If electrons are only accelerated by the reconnection electric field of the turbulent magnetic field, the acceleration timescale is estimated to be $t_{\rm acc}=\eta_{\rm acc}\gamma m_{e}c/(e\delta B)$. We estimate $\delta B$ as $\delta B\approx(\delta B/B)B_{\rm dis}$, and thus, the acceleration timescale is longer than that given in the text by a factor of $B/\delta B$. This affects the estimate of $\gamma_{\rm rad}$ but does not affect our results within the parameters of our interest. where $\eta_{\rm acc}\sim 10$ is the numerical factor and $e$ is the electric charge (e.g., Guo et al., 2020). We estimate $B_{\rm dis}$ by Equation (6) as
@@
We consider the cooling processes of the electrons as synchrotron radiation, synchrotron self-Compton scattering, and adiabatic cooling. As for the escape process, we consider that the electrons escape from the jet diffusely through resonant scattering with the turbulence induced by KHI. We explain how to estimate the cooling and the escape timescales in Appendix D.
@@ 3 Results for strong plasma entrainment
In this section, we apply the Jet-MAD model to M87 and compare the multi-wavelength photon spectra obtained by the Jet-MAD model to the observational data. We tabulate the parameters and the physical quantities for M87 in Table 1. We show the various timescales of the MAD model as a function of the proton energy $E_{p}$ for M87 in Figure 2. Due to the hard-sphere-like diffusion coefficient, the diffusion and acceleration timescales of the MAD model are independent of $E_{p}$. The cooling timescales are negligibly small for $E_{p}\lesssim 10^{9}~{}\rm GeV$.
We estimate the index of the injection term for the MAD, $s_{\rm inj}$, by the stochastic acceleration process with steady-state and hard-sphere approximations. As long as the cooling timescale is negligible, this process results in $s_{\rm inj}=-(1/2)+\sqrt{(9/4)+\epsilon}$, where $\epsilon=t_{\rm acc}/t_{\rm diff}$ is the escape-acceleration timescale ratio (see Section 2 of Stawarz & Petrosian, 2008). We obtain $s_{\rm inj}\simeq 1.1$ in our fiducial parameters.
@@
Parameters of the Jet-MAD model for strong plasma entrainment Parameter Description $\dot{m}$ $6.0\times 10^{-5}$ Mass accretion rate normalized by $L_{\rm Edd}/c^{2}$ $M_{\rm BH}~{}[M_{\odot}]$ $6.3\times 10^{9}$ BH mass $r~{}[R_{g}]$ 10 Size of the emission region for the MAD $Q_{i}/Q_{e}$ 10 Ratio of the electron-proton heating rate $\eta_{\rm turb}$ 0.833 Numerical factor of the effective mean free path $\alpha$ 0.3 Viscous parameter $\beta$ 0.1 Plasma beta $\epsilon_{\rm NT}$ 0.33 Fraction of the nonthermal particle energy to the dissipation energy $\epsilon_{\rm dis}$ 0.15 Fraction of the dissipation energy to the accretion energy $\sigma_{\rm ent}$ 0.01 Magnetization parameter after plasma entrainment at the jet dissipation region $p$ 2.16 Index of the injection term for the jet $\xi$ $10^{3}$ Numerical factor of the blob expansion $\delta B/B$ 0.35 Ratio of the amplitude for the perturbed and the ordered magnetic fields inside the jet $R_{j}~{}[R_{g}]$ $10^{2}$ Lateral size of the jet dissipation region
Derived physical quantities of M87 Physical quantity Description $s_{\rm inj}$ 1.1 Index of the injection term for the MAD $\sigma_{\pm}$ $1.0\times 10^{5}$ Magnetization parameter at the jet base $B_{\rm dis}\rm[G]$ $6.5\times 10^{-2}$ Magnetic field strength at the jet dissipation region $n_{e}\rm\ [cm^{-3}]$ 22 Electron number density at the jet dissipation region $L_{j}\ \rm[erg~{}s^{-1}]$ $2.8\times 10^{43}$ Injected power at the jet base $L_{e}\ \rm[erg~{}s^{-1}]$ $1.7\times 10^{41}$ Electron luminosity at the jet dissipation region $\delta_{D}$ 1.9 Doppler factor
Figure 3: Photon spectrum for M87. The thick and thin lines are the photon spectra after and before internal attenuation by the Breit-Wheeler process, respectively. Gray points are the multi-wavelength simultaneous data taken from EHT MWL Science Working Group et al. (2021) and MAGIC Collaboration et al. (2020). Blue and orange shaded regions are power-law and log-parabola fitting spectrum taken from the Fermi-4FGL-DR3 catalog (https://fermi.gsfc.nasa.gov/ssc/data/access/lat/12yr_catalog/).
Figure 4: Same as Figure 2, but timescales in the jet as a function of the electron energy. $t_{\rm SSC}$ (dotted) and $t_{\rm ad}$ (dot-dashed) are the synchrotron self-Compton and adiabatic cooling timescales, respectively. The vertical dashed gray line represents $E_{e,\rm cut}=\gamma_{\rm cut}m_{e}c^{2}$.
First, we consider a case with $(\delta B/B)^{2}\sigma_{\rm ent}<1$. We show the results in Figures 3 and 4. Thermal synchrotron radiation reproduces the mm/sub-mm data. Synchrotron radiation by the nonthermal protons and secondary electron-positron pairs produced via the Bethe-Heitler process in the MAD reproduce the GeV and TeV gamma-ray data, respectively. Synchrotron radiation by the jet contributes to the optical, UV, and X-ray data. Our model can reproduce multi-wavelength data except for tens of GHz radio data. In our one-zone jet emission model, the optical depth of the synchrotron self-absorption at the GHz band is $\sim 1$. Higher-frequency radio data should be produced by more inner parts of the jet where the optical depth of the synchrotron self-absorption is higher. Thus, if we extend our one-zone model to a one-dimensional one to involve the emission from the inner parts of the jet, our model will be able to explain the radio data.
@@
×$\times\xi^{2}_{3}\sigma^{2}_{\pm,4}\left(\frac{\sigma_{\rm dis}/(1+0.5\sigma_{\rm dis})}{0.66}\right)^{1/2}\ \rm eV,$ (12)
where $h$ is the Planck constant. Thus, the electrons accelerated by the magnetic reconnection have a Lorentz factor high enough to emit the optical to the X-ray photons via synchrotron radiation. On the other hand, the photon flux of the synchrotron self-Compton scattering is much lower than the synchrotron flux as shown in Figure 3. In our model, synchrotron flux is calibrated by the optical data to be $F_{\rm opt}\simeq 8\times 10^{-13}\ \rm erg~{}s^{-1}~{}cm^{-2}$, and we estimate the photon energy density as $U_{\rm rad}=F_{\rm opt}D^{2}/(R_{j}^{2}c)\simeq 8.4\times 10^{-6}\ \rm erg~{}cm^{-3}$, where $D=17~{}\rm Mpc$ is the distance to M87. We also estimate the magnetic energy density to be $U_{B}=B_{\rm dis}^{2}/(8\pi)\simeq 1.7\times 10^{-4}\ \rm erg~{}cm^{-3}$. The energy density ratio is estimated to be $U_{\rm rad}/U_{B}\simeq 0.05$, and thus the photon flux of the synchrotron self-Compton scattering is much lower than the synchrotron component.
Figure 4 shows the various timescales in the jet as a function of the electron energy. For our fiducial parameters, electrons are efficiently scattered by the turbulence, and the diffusive escape is inefficient. The nonthermal electrons are mainly cooled by the adiabatic expansion and the synchrotron radiation for $\gamma<1.9\times 10^{4}$ and $\gamma>1.9\times 10^{4}$, respectively. The photon energy emitted by the electrons with $t_{\rm ad}=t_{\rm syn}$ is estimated to be $E_{\gamma,\rm cool}\simeq 0.29~{}\rm eV$, at which the synchrotron spectrum is the peak, and that enables our model to explain the optical data.
@@ 4 Resutlts for weak or no plasma entrainment
The obtained minimum Lorentz factor is too high to emit the optical and X-ray photons. Based on Equation (13), $\delta B/B\sim 0.01$ is required to explain the X-ray data. Even if we tune $\delta B/B$ to emit the X-ray photons by electrons with $\gamma_{\rm min}$, the photon spectrum is too peaky to fit the optical and the X-ray data simultaneously. For the case of weak plasma entrainment, we show the result in Figure 5 and tabulate the changed parameters and physical quantities in Table 2. Other parameters and physical quantities are the same as those tabulated in Table 1. We conclude that the weak plasma entrainment is disfavored to explain the multi-wavelength data.
Table 2: The list of the changed parameters and the physical quantities for the case of the weak plasma entrainment.
$\sigma_{\rm ent}$ $\sigma_{\rm dis}$ $\delta B/B$ $B_{\rm dis}~{}[\rm{G}]$ $n_{e}~{}\rm[cm^{-3}]$ $L_{e}~{}\rm[erg~{}s^{-1}]$
Thick $5.5\times 10^{4}$ $9.2\times 10^{2}$ $0.033$ $0.92$ $8.2\times 10^{-4}$ $6.4\times 10^{40}$
Thin $5.5\times 10^{4}$ $2.5\times 10^{3}$ $0.02$ $0.92$ $8.2\times 10^{-4}$ $2.4\times 10^{40}$
Figure 5: Same as Figure 3, but for the weak plasma entrainment. Total and jet synchrotron photon spectra with $\delta B/B=0.033$ (thick lines) and $\delta B/B=0.02$ (thin lines). We set $\delta B/B=0.02$ to emit the optical photons and $\delta B/B=0.033$ to emit the X-ray photons. Both photon spectra cannot reproduce the multi-wavelength data simultaneously. Due to high $\sigma_{\rm dis}$, the photon flux of the synchrotron self-Compton scattering is low.
@@ 5.1.1 Power-law tail in the jet dissipation
We apply the Jet-MAD model with the power-law tail distribution to M87 and find that the Jet-MAD model can explain the multi-wavelength data even if the differential electron number spectrum has the power-law tail. The index of synchrotron flux is $s\approx(2-p)/2$ for the fast cooling regime. We obtain $s=-0.25$ and $-0.5$ for $p=2.5$ and 3.0, respectively, with which the synchrotron flux emitted by the electrons in the tail ($\gamma>\gamma_{\rm cut}$) has little effect on the overall photon spectrum. Even if the tail index is a bit hard, the photon energy higher than $E_{\gamma,\rm max}$ does not affect the photon spectrum (see Figure 3). We conclude that our model can explain the multi-wavelength data independent of the uncertainty for the reconnection acceleration.
@@ 5.1.2 Power-law tail for the electron-positron pair injection into the jet
Next, we evaluate the effect of the power-law tail on the electron-positron pair injection at the jet base. With the power-law tail, the photon spectrum extends up to a higher energy than that without the tail. One might expect a high pair production efficiency, compared to that estimated with the exponential cutoff as in Chen et al. (2023). We estimate the pair production rate and magnetization parameter at the jet base with the power-law tail distribution (see Appendix E for details). We find that the estimated magnetization parameters with the tail are comparable to that in our fiducial treatment: $\sigma_{\pm}=8.8\times 10^{4}$ and $1.0\times 10^{5}$ for $p_{\rm tail}=2.5$ and $3.0$, respectively. We show the estimated magnetization parameter as a function of the tail index in Figure 6. As shown in Figure 6, the estimated magnetization parameter hardly changes its value with the tail index. This comes from the self-regulation of the synchrotron spectrum to peak at $E_{\gamma,\rm cut,ms}\sim m_{e}c^{2}$. Therefore, the power-law tail does not affect our conclusions.
@@ 5.2 Acceleration by the Alfvén wave dissipation
In this subsection, we discuss the Alfvén waves dissipation as an alternative particle acceleration mechanism in the jet (Nättilä & Beloborodov, 2022). Alfvén waves are produced by the magnetic reconnection in the BH magnetosphere or by the interaction of the jet with disk wind. These Alfvén waves propagate along the magnetic field lines and split into forward sound waves and backward Alfvén waves by the parametric decay instability (PDI) (Galeev & Oraevskii, 1963; Sagdeev & Galeev, 1969; Derby, 1978). We consider that the Alfvén wave turbulence is generated by the interaction of the forward Alfvén waves with backward Alfvén waves at the dissipation region, and the dissipation of Alfvén wave turbulence accelerates the electrons. It is shown that in a highly magnetized medium, the PDI can split the Alfvén waves into backward Alfvén waves and forward sound waves only when $(\delta B/B)^{2}\sigma\leq 1$ is satisfied (Ishizaki & Ioka, 2024)  [^3] If the Alfvén waves are in $(\delta B/B)^{2}\sigma\geq 1$, they can split into backward Alfvén and backward sound waves (Ishizaki, private communication). In this case, due to the high growth rate of the PDI, the dissipation region is much smaller than that we consider in this paper, resulting in the high optical depth of the synchrotron self-absorption. If we tune $\sigma_{\rm ent}$ to avoid the high optical depth of the synchrotron self-absorption, our model with the Alfvén wave dissipation scenario may explain the multi-wavelength data.. We set $\delta B/B=1/\sqrt{\sigma_{\pm}}\simeq 3\times 10^{-3}$ which is close to the maximum value of $\delta B/B$ that satisfies the PDI growth condition. Then, the maximum Lorentz factor is written as $\gamma_{\rm max}\approx\xi$ by Equation (9), and the value is lower than that of the magnetic reconnection scenario.
Figure 8: Same as Figure 4, but for the acceleration by the Alfvén wave dissipation.
We show the results in Figures 7 and 8. We estimate the synchrotron photon energy emitted by electrons with $\gamma_{\rm max}$ to be
@@
Diffusive escape also affects the photon spectrum. For the Alfvén wave dissipation scenario, the amplitude of the turbulent magnetic field is lower than that of the magnetic reconnection scenario. Due to the low amplitude, the pitch angle scattering rate is low, leading to a high spatial diffusion coefficient. The diffusive escape timescale is always shorter than the adiabatic and synchrotron cooling timescales, causing a low radiation efficiency, compared to the magnetic reconnection scenario (see Figure 8). As a result, the calculated photon flux is low for the Alfvén wave dissipation scenario. We conclude that the dissipation of the Alfvén wave turbulence is disfavored as the acceleration mechanism inside the jet.